# Label Anything: Multi-Class Few-Shot Semantic Segmentation with Visual Prompts

### v1

@@ 3.1 Problem Formulation
Figure 2: Class prototype generation process. LA begins with extracting image features from a single support set example using ViT. This example might include multiple prompts related to different classes and types. A CNN-based Mask Encoder processes dense prompts (masks), while point coordinates and bounding boxes are transformed into embeddings via the Point Encoder, with bounding boxes treated as pairs of point embeddings. These processed prompts are then integrated with the image features. The Class Token Pool enriches dense and sparse embeddings with class information to infuse class-specific semantics. Subsequent two-way cross-attention between dense and sparse features yields class prototypes for each example.
3.2 Model Architecture
Label Anything processes a query image $I_{q}$ for segmentation alongside a support set $\mathcal{S}$, utilizing three core components in its architecture (refer to Fig. 1): the Image Encoder, the Prompt Encoder, and the Mask Decoder.
The Image Encoder leverages a pre-trained Vision Transformer, specifically ViT-B/16 [5] to process the query and support images. These images are transformed into feature maps $F_{q}$ for the query and $F_{s}$ for the support set, with dimension $D_{ViT}\times H_{d}\times W_{d}$, where $H_{d}=H/x_{p}$ and $W_{d}=W/x_{p}$ and $x_{p}$ is the ViT patch size. Attached to ViT is a convolutional neck that processes the image features to reduce their dimensionality from $D_{ViT}=768$ to $D=512$.
The Prompt Encoder is a key component of the LA architecture, designed to embody different types of prompts—masks, boxes, and points—from support images and effectively encapsulate this information. In its initial stage, the Prompt Encoder processes these inputs to generate embeddings: masks are encoded using a CNN, resulting in dense embeddings ($P_{d}$), while points and boxes are encoded through a Point Encoder, which uses Positional Embeddings to produce sparse embeddings for points and boxes ($P_{s}$). Three different learnable embeddings are used, each corresponding to the three types of coordinate inputs (point prompt, box top-left corner, box bottom-right corner), which are then summed. To enable attention on various objects as dictated by sparse prompt features, a self-attention layer complemented by a residual connection is employed. The layer operates on an input dimensionality of $BL\times NM\times D$, signifying batch size ($B$), number of examples ($L$), classes ($N$), prompts per class ($M$), and feature size ($D$), thereby facilitating independent example processing and attention to each prompt feature across both similar and different classes. We adopt the Token Pool strategy from Zhang et al. [41] to incorporate class-specific information using a trainable random Gaussian matrix sized $R\times D$, where $R$ should be greater than the number of classes. At each step, a row is sampled from this matrix for each class and added to the sparse and dense features. Subsequently, dense embeddings are combined with the image features to highlight the featured object within the feature map. Merging sparse and dense embeddings in a single feature map requires a Two-Way Transformer, similar to the mechanism in SAM [9]. This Transformer alternates attention between dense to sparse features and vice versa, according to:
$P_{s}$ $=MultiHeadAttention(P_{s},P_{d}+PE,P_{d}+PE),$ (1)
$P_{d}$ $=MultiHeadAttention(P_{d}+PE,P_{s},P_{s}),$ (2)
where $PE$ denotes the positional encoding adjusted for dimensions $(H,W)$. This process ensures a comprehensive integration of prompt-based information, highlighting the relevance of each object within the feature map.
After processing sparse features, these features are eliminated, forcing the model to encapsulate all pertinent details within the dense features. A Global Average Pooling layer then condenses the spatial dimensions, generating a set of class-example embeddings $E=\{E_{ni}\in\mathbb{R}^{D}|\forall n\in\mathcal{C},\forall i\in\mathcal{S}\}$. To facilitate the interchange of information across examples and classes, we introduce the Class Example Mixer block, which enables communication between different examples through a Transformer layer. The attention layer in it enables the embeddings of the same class $n$ over different examples to be attended:
$E_{n}$ $=E_{n_{1}},E_{n_{2}},\dots,E_{n_{L}}$ (3)
$E_{n}$ $=MultiHeadAttention(E_{n},E_{n},E_{n})$ (4)
Embeddings are averaged across examples to yield a singular prototype per class, compacting essential class information. Figure 2 shows a detailed view of the class prototype generation for a single support set example.
The Mask Decoder uses the class prototypes $E_{1},\dots,E_{|\mathcal{C}|}$ (or class-example prototypes at test time), and the query image features $F_{q}$ to generate a segmentation mask for each class. This process begins with the Two-Way Transformer, which serves a role different from its Prompt Encoder function. Here, it aims to facilitate pattern matching between class prototypes and pixel locations within the query image by reciprocally attending to them, thus transferring relevant information for segmentation. Subsequently, the query features are upscaled to a resolution of $(H/4,W/4)$, and their dimensionality is reduced to $D/8$ through a sequence of transposed convolutional layers. Concurrently, the class prototypes are dimensionally aligned to match the query features via an MLP, ensuring uniformity in representation. Following this, the modified query features $F_{q}$ undergo further refinement through three $3\times 3$ spatial convolutional layers aimed at producing segmentation outputs that are not only smoother but also spatially coherent. The segmentation logits for each pixel and class are determined through a dot product operation, succinctly captured by the formula:
$\text{Logit}_{xy,n}=F_{q_{xy}}\cdot E_{n},$ (5)
thereby computing the final segmentation masks.
@@ 3.3 Training Procedure
One of the primary objectives of this work is to enable our model to be trained once for application across arbitrary combinations of $N$ and $K$ during testing and across various prompt inputs. To accomplish this, we introduce an innovative episodic training approach that optimizes resource utilization while accommodating diverse $N$, $K$, and prompt-type configurations.
The procedure starts by defining potential batch configurations as $(B,N,K)$, where $B$ represents the batch size, indicating the number of learning episodes per batch. The selection of $B$ is tailored to the $N$ and $K$ parameters to ensure optimal GPU resource management using powers of 2, with $B\times N\times K+B$ representing the aggregate image count per batch, including both support and query images. Each training batch is then constructed by randomly selecting a $(B,N,K)$ tuple.
For every episode within a batch, $N$ classes are randomly chosen from the set of seen classes, $\mathcal{C}_{seen}$. A coin flip probabilistic decision determines whether the query image should contain any or all of the $N$ selected classes, adjusting $N$ if a comprehensive query image is unavailable. This variability fosters episodes where the query image encompasses objects from all sampled classes and scenarios, necessitating the exclusion of certain support set classes during segmentation, thereby addressing a recognized bias in FSS research [8].
Subsequently, $K$ support examples per class are sampled. We identify all annotations corresponding to the selected $N$ classes within each chosen image, allowing multiple annotations per image. This flexibility ensures our model’s readiness for any $L$ image-annotation pair configuration, where $L$ may deviate from the strict $N\times K$ formula.
Additionally, our training episodes are enriched by incorporating a random selection of prompt types for each instance annotation within the support set, enhancing the diversity of learning stimuli. For instances requiring point annotations, points are sampled proportionately to the instance’s mask area, tailoring the learning process to the specific needs of each episode and prompt type.
@@ 4.1 Dataset and Validation
Our assessment was based on the COCO-$20^{i}$ dataset [22, 33], a predominant FSS benchmark sourced from the MS COCO dataset [12], to evaluate our model across both $1$-way $1$-shot and $N$-way $K$-shot segmentation tasks. COCO-$20^{i}$ amplifies the challenge by dividing its 80 classes into four distinct folds, each encompassing 20 classes. Employing cross-validation, we rotated classes within each fold as $\mathcal{C}_{unseen}$, contrasting with $\mathcal{C}_{seen}$ for the remaining classes. This setup aligned $\mathcal{D}_{train}$ and $\mathcal{D}_{test}$ with the original dataset’s arrangement, ensuring our model faced novel images and classes during testing. Our model was fine-tuned in a single iteration per cross-validation phase, highlighting our training strategy’s effectiveness.
To ensure a rigorous comparison with established benchmarks, especially in binary FSS contexts [21, 22], we adopted a consistent evaluation methodology, conducting tests over $1000$ randomly chosen episodes, utilizing $5$ distinct random seeds for each and averaging the outcomes. This procedure was replicated for $N$-way $K$-shot evaluations, aligning with methodologies from prominent studies [42, 33, 14]. Unless otherwise stated, we limited our prompt type to masks during tests for more accurate benchmarking, facilitating direct comparisons across different approaches and configurations.
@@ 4.2 Setting
We adopted a uniform input resolution of $1024\times 1024$ for query and support images, resizing images and applying padding to maintain the aspect ratio. Following the methodology proposed by Zhang et al. [41], we used the ViT-B/16 model [5], augmented with Segment Anything [9] pretraining and with its weights kept frozen, serving as our backbone. This configuration yielded a singular feature map per image. Class-specific masks were supplied at a refined resolution of 256 $\times$ 256, while points and bounding boxes were integrated into the model using normalized coordinates.
Moreover, for a fairer comparison with existing methods, we also employed the same ViT-B/16 architecture with the Masked Autoencoder (MAE) pre-training strategy [2]. The input size was $480\times 480$, with the output feature map of dimension $768\times 30\times 30$. This pretraining strategy is preferred against the classical ImageNet classifiation pre-training as it retains the spatial features useful for segmentation.
Our training strategy optimized the focal loss [11], which addresses class imbalance by focusing on difficult-to-classify examples computed for a single training instance as:
$\mathcal{L}=\frac{1}{N+1}\sum_{n}^{N+1}{\left[w_{n}\cdot(1-e^{-l_{ce}(\hat{y}_{n},y_{n})})^{\gamma}\cdot l_{ce}(\hat{y}_{n},y_{n})\right]},$ (6)
where $N+1$ is the number of classes, $w_{n}$ are class-specific weights, $l_{ce}$ the cross-entropy loss, and $\gamma$ the focusing parameter adjusting the emphasis on hard examples.
We used the AdamW optimizer [19] ($\beta_{1}=0.9$, $\beta_{2}=0.999$), complemented by a linear learning rate warmup [7] for 1000 iterations, followed by a step-wise cosine learning rate decay schedule [18]. The initial learning rate, after warmup, was $1\mathrm{e}{-5}$. Our training followed the configurations outlined in Sec. 3.3, adopting batch size-number of classes-number of shots configurations such as $(4,1,4)$, $(2,4,2)$, $(8,1,2)$, $(4,2,2)$, $(4,4,1)$, and $(16,1,1)$, where the batch size refers to a single GPU. The model underwent training for 50 epochs, with the iteration count per epoch equivalent to the size of $\mathcal{D}_{train}$. A cap was placed on point samples at a maximum of $10$ per instance to maintain efficiency. The architecture totals 27 million learnable parameters, reflecting the model’s complexity and capacity for handling diverse FSS scenarios.
Training leveraged the Leonardo cluster’s resources, including 512GB of RAM and four NVIDIA A100-64GB GPUs, to support the computational demands, ensuring high-performance processing capabilities.
@@
fold-0 fold-1 fold-2 fold-3 mean fold-0 fold-1 fold-2 fold-3 mean
DCAMA [27] 49.5 52.7 52.8 48.7 50.9 55.4 60.3 59.9 57.5 58.3
FPTrans [40] 39.7 44.1 44.4 39.7 42.0 49.9 56.5 55.4 53.2 53.8
LA (MAE) 35.2 44.0 39.4 40.4 39.7 38.5 44.8 39.5 42.3 41.2
LA (SAM) 39.2 46.6 44.1 42.7 43.1 42.4 49.3 45.9 43.4 45.1
Table 1: Results for $1$-way $K$-shot segmentation, benchmarking against leading transformer-based FSS methodologies. DCAMA [27] represents the peak in affinity learning-based segmentation, leveraging the hierarchical Vision Transformer Swin-B [16] as its foundation. Conversely, FPTrans [41] exemplifies top-tier performance in prototype-based segmentation, employing the ViT-B/16 [5] as its backbone.
4.3 Results
In Table 1, we showcase the performance of LA on the COCO-$20^{i}$ benchmark, juxtaposing it with leading binary FSS methods. Our model achieved state-of-the-art results among prototype-based approaches, notably surpassing FPTrans [41] in the 1-shot setting. However, DCAMA [27], an affinity-based method, exceeds our performance by comparing each pixel in the query image against every pixel in the support images, leveraging features from multiple backbone levels. Given its detailed pixel-wise comparison strategy, DCAMA’s superior performance was anticipated. Despite this, affinity-based methods face scalability challenges, particularly in $N$-way $K$-shot scenarios, as the number of necessary comparisons increases quadratically with the image count and, by extension, with the class count. This limitation underscores the practical advantages of prototype-based models like LA for complex segmentation tasks involving multiple classes and shots.
Indeed, Table 2 presents the performance of LA and several other state-of-the-art methods on the $2$-way $K$-shot segmentation task, as evaluated on the COCO-$20^{i}$ benchmark. LA demonstrates superior performance across all four folds, excelling in the more challenging $2$-way $1$-shot and $2$-way $5$-shot scenarios, achieving the highest mIoU scores. Notably, LA surpassed MFNet [42], previously regarded as the leading method, especially in folds 1, 2, and 3, underscoring its superior ability to handle multi-class segmentation tasks with limited examples. Other methods show competitive but lower performance compared to LA, highlighting the effectiveness of our approach in leveraging diverse visual prompts and an end-to-end training strategy.
Furthermore, to validate the performances of LA in complex FSS settings, we tested our model capabilities also on an increasing $N$ setting. Table 3 shows the results of LA and DCAMA on the $N$-way $1$-shot segmentation task. Except for 1-way, LA outperforms DCAMA in all other $N$-way settings, showcasing its robustness and adaptability with a growing number of classes.
Figure 3 presents qualitative results from our $2$-way $1$-shot segmentation experiment on the COCO-$20^{i}$ dataset. This visual representation highlights the efficacy of our model in accurately interpreting and responding to diverse prompts, showcasing its robust segmentation capabilities in complex scenarios.
@@ 4.3 Results
LA (MAE) 30.6 33.6 32.2 31.2 31.9 21.1 25.1 26.4 27.5 25.1
LA (SAM) 31.0 38.2 34.7 34.5 34.6 24.2 27.5 29.6 29.7 27.7
Table 2: Performance on $2$-way / $5$-way $1$-shot segmentation benchmarked with contemporaneous methods as re-evaluated in the MFNet study [42]. Comparisons are made using mIoU percentages, specifically focusing on foreground classes. Misclassifications within these classes are considered under false positive and false negative categories for the respective classes; a measurement approach denoted as mIoU* in the MFNet analysis.
Method 1-way 2-way 3-way 4-way 5-way 10-way 15-way 20-way
DCAMA 50.9 31.7 24.2 21.1 16.7 10.8 6.5 4.7
LA (MAE) 39.7 31.9 28.6 26.5 25.1 20.7 15.0 10.8
LA (SAM) 43.1 34.6 31.7 29.6 27.7 23.6 16.9 13.7
Table 3: Average mIoU results for $N$-way 1-shot on COCO-$20^{i}$.
Figure 3: Visual representation of $2$-way $1$-shot segmentation on COCO-$20^{i}$. Sequence from left to right includes the first prompt, the second prompt, the ground truth mask, and our model’s prediction.

### v2

@@ 4 Proposed Method
Figure 2: Class prototype generation. LA extracts image features from a single support example via ViT. Dense prompts (masks) are encoded with a CNN; point coordinates and boxes are embedded via a point encoder, with boxes as point pairs. Prompt embeddings are fused with image features. A class token pool enriches them with class semantics. Two-way cross-attention between dense and sparse features produces the class prototypes. Class Example Mixer allows class information exchange.
4.1 Feature Extraction and Neck
We extract features from both the query image and the support set using a Vision Transformer, specifically ViT-B/16 [5], initialized with ImageNet pretraining in line with common practice in FSS [39]. By employing a ViT, we avoid aggregating multi-level feature maps from different layers, as is typically required with CNN-based encoders. Instead, we use the output from the final layer of the ViT, resulting in feature maps denoted as $F_{q}$ for the query and $F_{\mathcal{S}}$ for the $L$ support images. These have dimensionality $D_{\text{ViT}}\times H_{d}\times W_{d}$, where $H_{d}=H/x_{p}$ and $W_{d}=W/x_{p}$, with $x_{p}$ representing the patch size (e.g., $16$ for ViT-B/16).
To reduce the channel dimensionality and facilitate subsequent processing, we append a convolutional neck to the ViT. This module reduces the feature dimension from $D_{\text{ViT}}=768$ to a more compact representation with $D=512$ channels, yielding the final image features used in the following stages of the model.
4.2 Prototype Learning
Inspired by the architecture of the Segment Anything Model [9], we design our model to handle multiple types of prompts—masks, bounding boxes, and points—applied to the support images to define the concepts to be segmented in the query image. To this end, we extend the prompt encoder mechanism from SAM to suit the requirements of few-shot semantic segmentation.
Prompt Encoding
We begin by extracting embeddings from the input prompts. Mask prompts are processed through a series of convolutional layers, producing dense embeddings denoted as $P_{d}$. Point and box prompts are handled by a point encoder, which applies positional encodings to the $(x,y)$ coordinates of the input points, resulting in sparse embeddings $P_{s}$. We introduce three learnable type-specific encodings corresponding to point prompts, box top-left corners, and box bottom-right corners to distinguish between different types of sparse prompts. These are added to the respective sparse embeddings. All prompt embeddings—dense and sparse—are projected to have a dimensionality of $D$, matching the channel dimension of the image feature maps, which facilitates subsequent integration within the network. Once these embeddings are obtained, we contextualize the sparse prompts by allowing them to exchange information via a self-attention layer followed by a residual connection. This attention mechanism operates independently on the prompts of each support image, without sharing information across different support examples.
A key distinction from SAM lies in the fact that, in our setting, prompts may correspond to different semantic classes that the user wishes to segment. To account for this, we introduce a token pool, implemented as a learnable Gaussian-initialized matrix of size $R\times D$, where $R$ is a large integer greater than the maximum number of classes $N$ used during testing. For each episode (i.e., a support-query pair), one row from this matrix is sampled per class and added to the corresponding dense and sparse embeddings. This mechanism provides the model with explicit information about which prompts belong to the same class, thereby improving class-level disentanglement and representation consistency during segmentation.
Image-Prompt Fusion
After extracting representations for the support images and their corresponding prompts, we aim to fuse this information to construct class prototypes that can effectively guide the segmentation of the query image.
To this end, we adopt a two-way transformer mechanism inspired by SAM, where multi-head attention is applied bidirectionally, from the sparse prompts to the image features and vice versa. Information exchange is restricted to each prompt and its associated support image. Before this fusion step, we incorporate the dense prompts (i.e., masks) into the image features via element-wise addition, yielding enriched image representations defined as
$F_{l,c}=F_{l}+P_{d}^{l,c},\quad\text{for }l=1,\dots,L,$ (1)
where each class $c$ is associated with a distinct feature map.
The fusion process is then carried out using the following multi-head attention layers:
$P_{s}^{l,c}$ $=\text{MultiHead}(P_{s}^{l,c},F_{l,c}+PE,F_{l,c}+PE),$ (2)
$F_{l,c}$ $=\text{MultiHead}(F_{l,c}+PE,P_{s}^{l,c},P_{s}^{l,c}),$ (3)
where $\text{MultiHead}(Q,K,V)$ denotes the standard multi-head attention operator with queries $Q$, keys $K$, and values $V$. Here, $P_{s}^{l,c}$ represents the sparse prompt embeddings for class $c$, and $PE$ denotes positional encodings adapted to the spatial dimensions $(H_{d},W_{d})$.
After fusion, we retain only the updated dense features $F_{l,c}$, which now encapsulate information from both the support image and the sparse prompts. We apply global average pooling over the spatial dimensions to obtain a compact representation, producing a class-example embedding $e_{l,c}\in\mathbb{R}^{D}$ for each support image and target class. Before aggregating these embeddings across support examples, we introduce the class-example mixer, a self-attention block designed to refine class representations. For each class $c$, we concatenate the embeddings of its support examples into a matrix $E_{c}=\left[\begin{smallmatrix}e_{1,c}&\dots&e_{L,c}\end{smallmatrix}\right]$, and apply the following operation:
$e_{c}=\text{MeanPool}(\text{SelfAttn}(E_{c})),$ (4)
where SelfAttn denotes a self-attention layer applied across the example dimension. The resulting output $e_{c}$ serves as the class prototype, capturing a rich, aggregated representation of the class to be segmented. These prototypes are subsequently used during the query image mask decoding process.
4.3 Mask Decoding
The mask decoder takes as input the class prototypes $E=[e_{c}]_{c=1}^{N}$ and the query image features $F_{q}$, and produces a segmentation mask for each target class. This process begins with a two-way transformer module, which serves a different purpose than in the prototype learner. Its goal is to facilitate pattern matching between the class prototypes $E$ and the pixel locations within the query image $F_{q}$ by enabling mutual attention. This reciprocal interaction transfers relevant class-specific information to the query features, guiding segmentation. Formally, this attention mechanism is defined as:
$F_{q}$ $=\text{MultiHead}(F_{q},E,E),$ (5)
$E$ $=\text{MultiHead}(E,F_{q},F_{q}).$ (6)
Next, the query features are upsampled to a spatial resolution of $(4H_{d},4W_{d})$, and their channel dimension is reduced to $D/8$ using a sequence of transposed convolutional layers. In parallel, the class prototypes are projected to the same feature dimension using an MLP, ensuring alignment with the transformed query features. The refined query features $F_{q}$ are then further processed through three spatial convolutional layers with kernel size $3\times 3$, which enhance the spatial coherence and smoothness of the segmentation output.
As a result, we obtain the final query representation $F_{q}\in\mathbb{R}^{{4H_{d}}\times 4W_{d}\times\frac{D}{8}}$ and the aligned class prototype matrix $E\in\mathbb{R}^{N\times\frac{D}{8}}$. The predicted segmentation mask is computed as follows:
$\hat{M}=F_{q}\cdot E^{\top},$ (7)
where each element $\hat{M}_{xyn}$ represents the logit for pixel $(x,y)$ to belong to the $n$-th class.
@@ 4.4 Training Procedure
One of the primary objectives of this work is to enable our model to be trained once for applications across arbitrary combinations of $N$ and $K$ during testing, as well as for various prompt inputs. To accomplish this, we introduce an innovative episodic training approach that optimizes resource utilization while accommodating diverse $N$, $K$, and prompt-type configurations.
The procedure leverages “episodic training” [30], having flexible number of classes and shots, reflecting a real world scenario. Specifically, we defined potential batch configurations as $(B,N,K)$, where $B$ represents the batch size, indicating the number of learning episodes per batch. The selection of $B$ is tailored to the $N$ and $K$ parameters to ensure optimal GPU resource management, utilizing powers of 2. Specifically, $B\times N\times K+B$ represents the aggregate image count per batch, encompassing both support and query images. Each training batch is then constructed by randomly selecting a $(B,N,K)$ tuple.
For each episode in a training batch, $N$ classes are randomly sampled from the set of seen classes, $\mathcal{C}_{\text{seen}}$. A query image is then selected such that, with uniform probability, it contains either a subset or all of the $N$ chosen classes. If no query image satisfies the desired class configuration, $N$ is adjusted accordingly. This strategy enables the construction of episodes where the query image does not necessarily include all the classes in the support set, thereby mitigating a well-known bias in FSS research [8]. Subsequently, $K$ support examples are sampled per class. We identify all annotations corresponding to the selected $N$ classes within each chosen image, allowing multiple annotations per image. This flexibility ensures our model’s readiness for any $L$ image-annotation pair configuration, where $L$ may deviate from the strict $N\times K$ formula.
To increase variability during training, each episode includes a random selection of prompt types for the instance annotations in the support set. For instances requiring point-based annotations, points are sampled in proportion to the area of the instance mask. This approach tailors supervision to the specific characteristics of each episode and prompt type, leading to more robust model learning.
@@ 5.1 Dataset and Validation
Our assessment was based on the COCO-$20^{i}$ dataset [21, 31], a prominent FSS benchmark sourced from the MS COCO dataset [11], to evaluate our model across both $1$-way $1$-shot and $N$-way $K$-shot segmentation tasks. COCO-$20^{i}$ amplifies the challenge by dividing its 80 classes into four distinct folds, each encompassing 20 classes. Employing cross-validation, we rotated classes within each fold as $\mathcal{C}_{\text{unseen}}$, contrasting with $\mathcal{C}_{\text{seen}}$ for the remaining classes. This setup aligned $\mathcal{D}_{\text{train}}$ and $\mathcal{D}_{\text{test}}$ with the original dataset’s arrangement, ensuring our model faced novel images and classes during testing. Our model was fine-tuned in a single iteration per cross-validation phase, demonstrating the effectiveness of our training strategy.
To ensure a rigorous comparison with established benchmarks, especially in binary FSS contexts [20, 21], we adopted a consistent evaluation methodology. We conducted tests on 1,000 randomly chosen episodes, using five distinct random seeds for each and averaging the results. This procedure was replicated for $N$-way $K$-shot evaluations, aligning with methodologies from prominent studies [14, 31, 40]. Unless otherwise stated, we limited our prompt type to masks during tests for more accurate benchmarking, facilitating direct comparisons across different approaches and configurations.
@@ 5.2 Setting
We adopted a uniform input resolution for query and support images by resizing them to $480\times 480$. We used the ViT-B/16 architecture as a backbone, pre-trained with the masked autoencoder (MAE) strategy [2] on ImageNet. Our training strategy optimized the focal loss [12], which mitigates class imbalance by emphasizing hard-to-classify examples. To further account for intra-image class imbalance, we applied per-class weighting proportional to the number of pixels belonging to each class in the query image. We used the AdamW optimizer [18] ($\beta_{1}=0.9$, $\beta_{2}=0.999$), complemented by a linear learning rate warmup [7] for 1000 iterations, followed by a step-wise cosine learning rate decay schedule [17]. The initial learning rate, after warmup, was $1\mathrm{e}{-5}$. Our training followed the configurations outlined in Sec. 4.4, adopting batch size-number of classes-number of shots configurations such as $(4,1,4)$, $(2,4,2)$, $(8,1,2)$, $(4,2,2)$, $(4,4,1)$, and $(16,1,1)$, where the batch size refers to a single GPU. The model underwent training for 50 epochs, with the iteration count per epoch equivalent to the size of $\mathcal{D}_{\text{train}}$. A cap was placed on point samples at a maximum of $10$ per instance to maintain efficiency. The architecture totals 8.9 million learnable parameters, reflecting the model’s complexity and capacity for handling diverse FSS scenarios. Training leveraged the Leonardo cluster’s resources, including 512GB of RAM and four NVIDIA A100-64GB GPUs.
5.3 Results
Table 1: Results for $1$-way $K$-shot FSS segmentation, benchmarking against leading multi-class FSS methodologies and FPTrans.
@@ 5.3 Results
PANet [31] 23.0 33.8
PPNet(w/o U) [14] 25.7 36.2
PPNet [14] 27.2 36.7
MFNet [40] 34.9 39.2
FPTrans [38] 42.0 53.8
LA 39.7 41.2
In Table 1, we report the performance of LA on the COCO-$20^{i}$ benchmark for binary classification, comparing it against other multi-way approaches. We also include FPTrans [39], a dedicated binary classification model, as a reference. LA achieves the best results among the multi-way methods and performs comparably to FPTrans in the 1-way 1-shot setting, indicating that the architectural extensions introduced in LA preserve their effectiveness even in binary scenarios. However, LA shows limitations in mIoU when using multiple shots for the target class, underperforming compared to FPTrans. This highlights a trade-off: LA’s flexibility comes at the cost of reduced specialization compared to methods tailored for 1-way settings. Nevertheless, like other prototype-based methods, LA underperforms compared to affinity-based approaches such as DCAMA [26], which leverages multi-scale features and performs dense pixel-wise comparisons—achieving, for example, a score of 50.9 in the 1-way 1-shot setting. However, affinity-based methods face scalability challenges in $N$-way $K$-shot scenarios, as the number of comparisons grows quadratically with the number of pixels. Specifically, if $N_{q}$ and $N_{s}$ denote the number of pixels in the query and support images, the number of pairwise comparisons is $N_{q}\times N_{s}$. This computational overhead highlights the practical advantages of prototype-based models, such as LA, for segmentation tasks involving multiple classes and shots.
We tested LA’s performance under more complex conditions in increasingly larger $N$-way settings to further evaluate it. We also adapted DCAMA to a multi-class sequential variant to mitigate its original quadratic complexity and to assess whether it can maintain performance in this more scalable setup. Figure 3 presents the results of LA and the adapted DCAMA on the $N$-way 1-shot segmentation task. Except for the 1-way case, LA consistently outperforms DCAMA across all other settings, underscoring its robustness and adaptability as the number of classes increases. To our knowledge, this is the first time an FSS method has been scaled beyond five classes in the multi-class setting. This highlights a current limitation in the field and a promising direction for future investigation into the scalability and generalization of few-shot segmentation methods.
Figure 3: Average mIoU results for $N$-way 1-shot on COCO-$20^{i}$.
Table 2: Profiling comparison between DCAMA and Label Anything in a 5-way setting (GB/s).
Model 1-shot 2-shot 5-shot 10-shot
DCAMA 6.72 / 0.22 12.19 / 0.30 28.00 / 0.50 55.00 / 1.02
LA 0.57 / 0.08 00.68 / 0.10 00.99 / 0.14 01.50 / 0.21
Table 3: Performance on FSS benchmarks in $2$-way/$5$-way $1$-shot setting with multi-class FSS methods as re-evaluated in the MFNet study [40]. Comparisons are made using mIoU percentages, specifically focusing on foreground classes. Misclassifications within these classes are considered under false positive and false negative categories for the respective classes, a measurement approach denoted as mIoU* in the MFNet analysis.
@@
LA 30.6 33.6 32.2 31.2 31.9 21.1 25.1 26.4 27.5 25.1
We also evaluated the impact of increasing the number of shots in a 5-way setting. As shown in Table 2, DCAMA exhibits a near-linear growth in memory usage, reaching 55 GB at 10 shots. In contrast, LA demonstrates a sublinear trend, requiring only 1.5 GB at the same shot count.
Table 3 presents the performance of LA and several other state-of-the-art methods on the $2$-way $K$-shot segmentation task. LA demonstrates superior performance across all four folds, excelling in the more challenging $2$-way $1$-shot and $2$-way $5$-shot scenarios, achieving the highest mIoU scores. Notably, LA surpasses MFNet [40], previously regarded as the leading method, especially in folds 1, 2, and 3, underscoring its superior ability to handle multi-class segmentation tasks with limited examples.
We further evaluated LA in a non-exclusive class setting, where training and testing classes are not disjoint, as explored in [33] (Generalist Models). For a fair comparison, LA was trained using features extracted from ViT-L, consistent with the backbone used by Painter [32] and SegGPT [33]. In this setting, LA demonstrates competitive performance, achieving state-of-the-art results in the 1-shot scenario and maintaining strong performance in the 5-shot scenario.
Figure 4: (Left) Visual representation of $2$-way $1$-shot segmentation on COCO-$20^{i}$. (Right) Visual representation of $2$-way, $L=3$ segmentation on COCO-$20^{i}$. The sequence from left to right includes the first prompt, the second prompt, the ground truth mask, and our model’s prediction.
Figure 4 (left) shows qualitative results for the standard $N=2$, $K=1$ segmentation setting on the COCO-$20^{i}$ dataset. On the right, we present segmentations for a more complex scenario with $N=3$ and $L=2$, where prompts for three classes are distributed across two support examples. This setup allows greater expressiveness and compositional freedom in defining segmentation tasks. Furthermore, this visual representation highlights the efficacy of our model in accurately interpreting and responding to diverse prompts, showcasing its robust segmentation capabilities in complex scenarios.
Table 4: Quantitative results on COCO-$20^{i}$ obtained by having categories in training covering categories in testing.
Method 1-way 1-shot 1-way 5-shot
Specialist Models
HSNet 41.7 50.7
VAT 42.9 49.4
FPTrans 56.5 65.5
LA 50.5 53.5
LA (L) 57.4 60.2
Generalist Models
Painter 32.8 32.6
SegGPT 56.1 67.9